OpenStudy (lexilove13):

Suppose a runner takes 45 min to run a route at 8mi/h at the beginning of training season. By the end of training season she can run the same route in 38 min. What is her speed at the end of training season?

OpenStudy (anonymous):

At the beginning of her season she was running at exactly 10.666666666666666 miles per hour. At the end of her training she was running at exactly 12.631578947368421.

OpenStudy (lexilove13):

How did you get that though @assaulth3ro911

OpenStudy (anonymous):

WEELLLL.... it all started one day... LOL jk (x Uhm well basically I realized they had no mention of the speed when she first started, so step one was to find the speed she was running at the beginning. Speed will be measured by MPH here, 45 minutes is only 3/4 of one hour. So in order to run 8 miles in one hour, you run at 8MPH, BUT, since she did it in 45 minutes it would essentially mean adding 25% (1/4) onto the speed. Now she has improved at the end of her training she has managed to run 8 miles in only 38 minutes (she's improving nicely, don't you think? (: ) So. You add to her speed again (because the time was reduced her speed clearly went up.) Now since I used a calculator to get this (LOL) I'm not sure how I got the final answer other than it's +2 again. Sorry I'm not a 100% genius at explaining this since I used a calculator and a bit of two sense. (; http://www.arkansasrunner.com/calculators/mileperhr.htm That is what I used. Maybe you can figure it out from that and what I said. (:

OpenStudy (kropot72):

\[45\ minutes=\frac{45}{60}\ hour=\frac{3}{4}\ hour\] Distance = time * speed \[\frac{3}{4}\times 8=6\ miles\] \[speed=\frac{distance}{time}\] \[38\ minutes=\frac{38}{60}\ hour\] \[speed=\frac{8}{\frac{38}{60}}=\frac{8\times 60}{38}=you\ can\ calculate\]

OpenStudy (anonymous):

What he said. ._. I knew I was right, I just had no clue on what I was saying. :D

OpenStudy (lexilove13):

Thank you both @assaulth3ro911 and @kropot72

OpenStudy (anonymous):

Anytime! ;D Have a wonderful day. (x

OpenStudy (kropot72):

You're welcome :)